# Improved Treatment of 1-4 interactions in Force Fields for Molecular Dynamics Simulations

### v2

@@ 4 Discussion and Conclusions
In contrast, the bonded-only model developed in this work circumvents these issues entirely for modeling 1-4 interactions. By replacing the hybrid treatment of torsional and non-bonded interactions with explicitly parameterized bonded coupling terms, we achieve exceptional accuracy in reproducing QM reference data across a diverse set of molecules, encompassing both flexible and rigid systems. The success of this approach highlights the importance of explicitly accounting for coupled bonded interactions, as we recently showed for bonds and angles27 and now extended to torsion-bond and torsion-angle, torsion-angle-angle terms, which have been underutilized in FF development in our view.
An additional and important feature of this work is the improved development of Q-Force25 for parameterizing these new bond coupling potentials. Notably, the errors in Q-Force predictions were even smaller than the discrepancies between the two QM methods used as benchmarks ($\omega$-B97X-V/def2-TZVPD and B3LYP-D3BJ/TZVP), suggesting that Q-Force closely approaches the inherent accuracy limits of the QM reference method. Importantly, the Q-Force toolkit is agnostic to the QM reference and the same methodology can be used to create FFs parameterized against coupled cluster, MP2 or any other QM method.25 Finally, the automated parameterization capabilities of Q-Force provide a reproducible and scalable framework for FF development in this regard as shown with ff14sb, CHARMM36, OPLS-AA, and is adaptable to advanced FFs by the same protocol as well.

### v3

@@ 4 Discussion and Conclusions
In contrast, the bonded-only model developed in this work circumvents these issues entirely for modeling 1-4 interactions. By replacing the hybrid treatment of torsional and non-bonded interactions with explicitly parameterized bonded coupling terms, we achieve exceptional accuracy in reproducing QM reference data across a diverse set of molecules, encompassing both flexible and rigid systems. The success of this approach highlights the importance of explicitly accounting for coupled bonded interactions, as we recently showed for bonds and anglesSami2024 and now extended to torsion-bond and torsion-angle, torsion-angle-angle terms, which have been underutilized in FF development in our view.
An additional and important feature of this work is the improved development of Q-ForceSami2021 for parameterizing these new bond coupling potentials. Notably, the errors in Q-Force predictions were even smaller than the discrepancies between the two QM methods used as benchmarks ($\omega$-B97X-V/def2-TZVPD and B3LYP-D3BJ/TZVP), suggesting that Q-Force closely approaches the inherent accuracy limits of the QM reference method. Importantly, the Q-Force toolkit is agnostic to the QM reference and the same methodology can be used to create FFs parameterized against coupled cluster, MP2 or any other QM method.Sami2021 Finally, the automated parameterization capabilities of Q-Force provide a reproducible and scalable framework for FF development in this regard as shown with ff14sb, CHARMM36, OPLS-AA, and is adaptable to advanced FFs by the same protocol as well.
@@
{suppinfo}
Supporting information: Error metrics, additional Ramachandran plots, and force field parameters.
{tocentry}[Uncaptioned image]